OpenStudy (anonymous):

what is the cos (17pi/4)

OpenStudy (anonymous):

cos ( 17Pi/4) = cos (16Pi/4 + Pi/4) = cos (4Pi + Pi/4) Use the trig identity: cos (a + b) = cos a*cos b - sin a* sin b cos 4Pi = cos 2Pi = 1; cos Pi/4 = V2/2 sin 4Pi = sin 0 = 0 cos(4pi + Pi/4) = ....